OpenStudy (anonymous):

If f(x) = 5x + 3 and g(x) = 2x + 3, what is f(g(x)) ?

OpenStudy (mathmate):

If f(x)=5x+2, and g(x)=3x+4 then f(g(x))=f(3x+4)=5(3x+4)+2=15x+20+2=15x+22 the given problem can be solved in a similar way.